OpenStudy (anonymous):

use l'hopital's rule to solve the limit as it approaches 0 from the right of 3/sinx - 1/x ??

OpenStudy (anonymous):

\[ \lim_{x \rightarrow 0+} 3/\sin x - 1/x\]

OpenStudy (anonymous):

limit is 0 l'hopitals rule you'll get lim of 0/cos(0) - 0/1 so 0

OpenStudy (anonymous):

so 0 is also an indeternminate form?

OpenStudy (anonymous):

thanks by the wayyyy!!! :)

OpenStudy (anonymous):

no, if the limit is 0, it exists it goes to 0

OpenStudy (anonymous):

if it comes out to be 0 / 0 then it's indeternminate

OpenStudy (anonymous):

First show that limit is indeterminate, then employ l'hopital's rule lim x->0 (3x-sinx)/(xsinx)

OpenStudy (anonymous):

i'm not sure if cyter has it right... i'm pretty sure you deriv instead of integral for l'hopital's rule

OpenStudy (anonymous):

this function ends up equaling 0.. thats where i am confused bcuz why do i need to use l'hopitals rule/.

OpenStudy (anonymous):

lim of 0/cos(0) - 0/1 ends up zero so the limit is zero but your original equation (the question) is undefined (or indeter.) so you have to apply l'hoptials' rule

OpenStudy (anonymous):

It's not 0. You can't use (and don't need) l'Hopital here.

OpenStudy (anonymous):

polpak, can you explain your reasoning?

OpenStudy (anonymous):

Err wait. It's not 0, but you do need l'hopital

OpenStudy (anonymous):

how is my original equation (the question) undefined? 3 over sin(0) - 1 over x = 3 over 0 - 1 over 0 = 0 - 0 = 0 ??

myininaya (myininaya):

polpak when you get a chance will you come back to that kid who is trying to teach himself

OpenStudy (anonymous):

3/0 is not 0 it's undefined. which is why i would do l'hopitals rule but polpak says i'm wrong :x so we'll see what he says first

OpenStudy (anonymous):

You don't have 0 - 0, you have \(3/0 - 1/0 = \infty - \infty\)

OpenStudy (anonymous):

And you do need lhopital, but you need to first put over a common denominator

OpenStudy (anonymous):

ohhhh *faceplam* i forgot limits >< it's "as you approach" not zero thanks polpak yup itenn, polpak's right

OpenStudy (anonymous):

\[3/sinx - 1/x = \frac{3x - sinx}{xsinx} \] Which gives a 0 over 0 and you take the derivative of top and bottom.

OpenStudy (anonymous):

Give that derivative a try and take the limit as the derivative aproaches 0

OpenStudy (anonymous):

I'll brb

OpenStudy (anonymous):

thankk you!!!!

OpenStudy (anonymous):

so far i have 3 - cos x all over 2 cos x - x sinx

OpenStudy (anonymous):

I agree with the numerator, but not the denominator.

OpenStudy (anonymous):

Using products rule \[\frac{d}{dx}x(sin\ x) = x(cos\ x) + (sin\ x)\]

OpenStudy (anonymous):

Which gives \[\frac{3-cosx}{x(cos\ x) + (sin\ x)}\] Which we can then plug in 0 for x

OpenStudy (anonymous):

2/0 ?

OpenStudy (anonymous):

Yep, \(2/0^+\)

OpenStudy (anonymous):

So \(+\infty\)

OpenStudy (anonymous):

why was the first equation indeterminate again.. im sorry im not getting this:(

OpenStudy (anonymous):

it was similat to the 2/0+ ... it was 3/0 - 1/0.

OpenStudy (anonymous):

Because you had 3/0 - 1/0. Which is \(\infty - \infty\)

OpenStudy (anonymous):

And that can be a lot of different things. It could be 0, it could be a constant, or it could be +/- infinity.

OpenStudy (anonymous):

So we have to put them over a common denominator to get 0/0 or \(\infty/\infty\) and then use l'Hopital's

OpenStudy (anonymous):

to see which infinity is growing faster

OpenStudy (anonymous):

ohkay. i did another on my own -- limit x approaches 1 [ x^3 - 1 all over x-1 ]

OpenStudy (anonymous):

i got 0/0 and then did derivative of top and bottom 3x^2 over 1

OpenStudy (anonymous):

and then i got 3. is that right?

OpenStudy (anonymous):

Yep, that's right.

OpenStudy (anonymous):

my last problem for l'hopitals rule is limit x approaches 0+ x^x ??

OpenStudy (anonymous):

what is the indeterminate form... if 0^0 = 0

OpenStudy (anonymous):

It's not 0. If you have \(\lim_{x \rightarrow c} f(x) = 0^0\) then you have to take the \[\lim_{e^x \rightarrow c} ln\ f(x)\]

OpenStudy (anonymous):

So in this case you have \[\lim_{e^x \rightarrow 0} x(ln\ x)\] Which is of the form \(0\cdot\infty\)

OpenStudy (anonymous):

0^0 is actually an indeterminate form so u can use l'hopital's rule

OpenStudy (anonymous):

and 0^0 is not equal to 0

OpenStudy (anonymous):

0 * inf is an indeterminate form?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

polpk, I have a quick question on mine you were helping me with.. when you're done here:)

OpenStudy (anonymous):

inf/inf, 0/0, -inf/inf, inf/-inf or 0^0, any of those are indeterminate forms

OpenStudy (anonymous):

\(0 * \infty\) is indeterminate.

OpenStudy (anonymous):

sometimes wikipedia is unreliable

OpenStudy (anonymous):

the next step then would be to take the derivativve?

OpenStudy (anonymous):

but after futhre digging I'm convinced

OpenStudy (anonymous):

Was gonna say, wolfram alpha also thinks it's indeterminate: http://mathworld.wolfram.com/Indeterminate.html

OpenStudy (anonymous):

Which next step? I may have a different process for this than cyter.

OpenStudy (anonymous):

im following you polpak..

OpenStudy (anonymous):

I'm not sure how he was gonna use l'hopital on 0^0 since there's no fraction there, but I'm a bit rusty with this

OpenStudy (anonymous):

thats where i confused.. l'hopitals rule has lim f(x)/g(x) and then f'(x)/g'(x)

OpenStudy (anonymous):

i dont see a fraction here.

OpenStudy (anonymous):

Since \[\lim_{e^x \rightarrow 0} x(ln\ x)\] is not nice, we have to swap our variable around a bit and say u = 1/x if x goes to 0, u goes to infinity. And that means that x = 1/u, so we have now \[\lim_{e^{u} \rightarrow \infty} \frac{ln\ 1/u}{u}\] Which gives us a nice fraction with \(\infty/\infty\) that we can take the derivative of.

OpenStudy (anonymous):

this one is very complicated.

OpenStudy (anonymous):

Yeah, a bit.

OpenStudy (anonymous):

i am not good with derivatices so the ln is trouble for me.. this is hard:(

OpenStudy (anonymous):

So \[\frac{d}{du} ln\ 1/u = u*(\frac{-1}{u^2}) = \frac{-1}{u}\]

OpenStudy (anonymous):

ohhh okay.

OpenStudy (anonymous):

so now we have Erg, my notation is wrong, but it's tiny if I fix it.

OpenStudy (anonymous):

what would change?

OpenStudy (anonymous):

Anyway, we have It's supposed to be e raised to the power of the limit of that whole thing, not e^u approaches infinity.

OpenStudy (anonymous):

So we have \[e^{\lim_{u \rightarrow \infty} \frac{-1}{t}} = e^0 = 1\]

OpenStudy (anonymous):

See, tiny.

OpenStudy (anonymous):

mm.. at which point?

OpenStudy (anonymous):

ohh it was never supposed to be e to the u

OpenStudy (anonymous):

Yeah, it was supposed to be e to the limit. That's it. The limit as \(x \rightarrow 0 \text{ of } x^x = 1\)

OpenStudy (anonymous):

so no l'hopital is needed??

OpenStudy (anonymous):

yes it is. I used it.

OpenStudy (anonymous):

This is too difficult with such a tiny font. Lemme try something else

OpenStudy (anonymous):

ohkay.

OpenStudy (anonymous):

i am confused:(

OpenStudy (anonymous):

im there..